# Tally: Non-Intrusive Performance Isolation for Concurrent Deep Learning Workloads

### v1

@@ 6. RELATED WORK
Many GPU sharing systems are designed with the goal of maximizing the throughput. Time-Slicing (GILMAN2021102234, ) and MPS (nvidia-mps, ) provided in NVIDIA GPUs enable transparent sharing of GPU resources in temporal and spatial manners. Besides vendor-provided options, many works (Ebird2019, ; gpupool2023, ; smgpu2016, ) have been built on the concept of co-executing kernels with complementary resource demands (e.g., compute vs. memory intensive) to further exploit resource parallelism in GPU SMs. For instance, Kernelet (zhong2013kernelet, ) and Slate (slate2019, ) perform such scheduling at runtime, while Kernel Fusion (fusion2010, ), HFUSE (li2022automatic, ), Tacker (tacker2022, ), and ISPA (ispa2023, ) employ source-to-source transformation to fuse kernels during compilation. Systems in this category fundamentally cannot support performance isolation as they prioritize the overall efficiency of the system over preservation of performance of certain high-priority workloads, thus falling short in production scenarios where tasks have SLAs.
Meanwhile, systems that are designed with performance isolation support, like Effisha (effisha2017, ) and GPES (gpes2015, ), utilize block-level preemption techniques to mitigate interference in concurrent execution but they require access to application source code. In contrast, Tally operates non-intrusively. AntMan (xiao2020, ) and TGS (tgs2023, ) leverage the iterative nature of DL training to adaptively limit the resources of low-priority tasks based on performance feedback from high-priority jobs. REEF (reef2022, ) implements a reset-based scheduling mechanism that achieves microsecond-scale kernel preemption latency on idempotent kernels. However, these works rely on specific task characteristics and cannot generalize to the wide variety of DL workloads. Conversely, Tally supports broad compatibility using task-agnostic scheduling and kernel transformations. Furthermore, Orion (orion2024, ) enhances performance isolation with kernel-level scheduling with consideration on each kernel’s compute and memory requirements. However, Orion requires offline profiling of workloads, which can be challenging to accomplish in cluster settings. Whereas Tally employs transparent profiling for priority-aware scheduling.
@@ 7. CONCLUSION
In this paper, we highlight the limitations of current GPU sharing systems and their inability to meet the versatile requirements of modern large-scale DL clusters. To address these limitations, we design Tally, a novel non-intrusive GPU sharing mechanism that offers robust performance isolation and universal applicability to diverse DL workloads. We empirically show that Tally can effectively ensure the performance standards of latency-critical tasks during shared execution, incurring a mere $7.2\%$ overhead on average on a diverse set of workload combinations, compared to 188.9% minimal overhead in state-of-the-art non-intrusive GPU sharing systems while offering competitive system throughput.

### v3

@@ 7. RELATED WORK
Many GPU sharing systems have been proposed with the goal of maximizing throughput. NVIDIA’s Time-Slicing (GILMAN2021102234, ) and MPS (nvidia-mps, ), for example, offer mechanisms to share GPU resources in temporal and spatial manners, respectively. Moreover, several studies (Ebird2019, ; gpupool2023, ; smgpu2016, ) have explored co-execution of kernels with complementary resource demands (e.g., compute vs. memory) to exploit resource heterogeneity in GPUs. Systems like Kernelet (zhong2013kernelet, ) and Slate (slate2019, ) schedule such co-execution dynamically at runtime, while approaches like Kernel Fusion (fusion2010, ), HFUSE (li2022automatic, ), Tacker (tacker2022, ), and ISPA (ispa2023, ) employ source-to-source transformations to fuse kernels during the compilation phase. Despite their efficacy in improving resource utilization, these systems inherently lack support for performance isolation, as they prioritize overall system efficiency over preservation of performance of individual workloads, making them unsuitable for production environments where tasks must adhere to strict SLAs.
To address the need for performance isolation, solutions like Effisha (effisha2017, ) and GPES (gpes2015, ) utilize block-level preemption techniques to mitigate interference during concurrent execution. However, these approaches require access to the application source code, whereas Tally operates non-intrusively. Other solutions, such as AntMan (xiao2020, ) and TGS (tgs2023, ), leverage the iterative nature of DL training to adaptively limit the resources of low-priority tasks based on performance feedback from high-priority jobs. Meanwhile, REEF (reef2022, ) implements a reset-based scheduling mechanism that achieves microsecond-scale kernel preemption latency on idempotent kernels. Yet, these works rely on specific task characteristics and cannot generalize to the wide variety of DL workloads. In contrast, Tally supports broad compatibility using task-agnostic scheduling and kernel transformations. Furthermore, Orion (orion2024, ) enhances performance isolation via kernel-level scheduling with consideration on each kernel’s compute and memory requirements. However, it requires offline profiling of workloads, which is challenging to accomplish in cluster settings. In comparison, Tally employs transparent profiling for priority-aware scheduling.
@@ 8. CONCLUSION
In this paper, we highlight the limitations of current GPU sharing systems and their inability to meet the diverse requirements of modern large-scale DL clusters. To overcome these challenges, we design Tally, a novel non-intrusive GPU sharing mechanism that offers robust performance isolation while maintaining compatibility with a wide range of DL workloads. Our evaluation shows that Tally can effectively ensure the performance standards of latency-critical tasks during shared execution, incurring a mere $7.2\%$ overhead on average while offering competitive system throughput.
Acknowledgements.
We thank the anonymous reviewers and our shepherd, Tim Rogers, for their valuable feedback. This research was supported in part by the Innovation JELF Grant, NSERC Discovery Grant, AWS Machine Learning Research Award (MLRA), Facebook Faculty Research Award, Google Scholar Research Award, and VMware Early Career Faculty Grant.
Appendix A Artifact Appendix
A.1. Abstract
We provide the source code of Tally along with scripts to reproduce experimental results presented in the evaluation section. This appendix includes instructions for generating plots similar to Figure 5 and Figure 7(b). Dockerfiles are provided to set up the required runtime environment. To expedite artifact evaluation, a pre-built image is available, containing the fully configured environment, pre-compiled deep learning models, and datasets as specified in Table 2. The experiments require an x86-64 Linux host with at least 85 GB RAM, 200 GB of free disk space, and an NVIDIA A100 GPU (40 GB). For consistent performance, we recommend using a Google Cloud a2-highgpu-1g instance.
A.2. Artifact check-list (meta-information)
  • Run-time environment: Docker.
  • Compilation: Pre-compiled within a Docker image.
  • Data set: MAF2 trace and datasets as listed in Table 2.
  • Hardware: A 12-core CPU with 85 GB RAM and an NVIDIA A100 GPU (40 GB).
  • Metrics: $99^{th}$ percentile latency and throughput.
  • Experiments: End-to-end evaluation (Section 5.2) and performance decomposition (Section 5.5). For faster evaluation, the default setup generates results only for co-executing BERT and Llama inference tasks alongside all training workloads. Instructions for running the full set of inference workloads can be found in Section A.7.
  • Output: Plots similar to Figure 5 and Figure 7(b).
  • How much disk space required (approximately)?: Approximately 200 GB.
  • How much time is needed to prepare workflow (approximately)?: About 10 minutes to download the Docker image.
  • How much time is needed to complete experiments (approximately)?: Approximately 6 hours for BERT and Llama tasks; 18 hours for the full end-to-end evaluation.
  • Publicly available?: Yes.
  • Code licenses (if publicly available)?: MIT License.
A.3. Description
A.3.1. How to access
The artifact is available for download on GitHub: https://github.com/tally-project/tally-bench.
A.3.2. Hardware dependencies
Requires an x86-64 Linux host with at least 85 GB of RAM, 200 GB of free disk space, and an NVIDIA A100 GPU (40 GB). We recommend a Google Cloud a2-highgpu-1g instance for consistent performance.
A.3.3. Software dependencies
Docker installation is required as all experiments are configured to run within the provided Docker image.
A.3.4. Data sets
The data sets used in the experiments are listed in Table 2 and are pre-downloaded as part of the Docker image.
A.3.5. Models
The models used in these experiments are listed in Table 2.
A.4. Installation
  • 1.
    Clone the Github repository.
    \$ git clone
      https://github.com/tally-project/tally-bench.git
    \$ cd tally-bench
  • 2.
    Pull the Docker image.
    \$ docker pull wzhao18/tally:bench
A.5. Experiment workflow
  • 1.
    Create the results directory and launch the Docker container.
    \$ mkdir tally_results
    \$ docker run -it --shm-size=64g --gpus 0 -v
      \${PWD}/tally_results:/home/tally-bench/tally_results
      wzhao18/tally:bench /bin/bash
  • 2.
    Run the following commands to run the experiments.
    (On host machine)
    \$ sudo nvidia-smi -i 0 -c DEFAULT
    (In docker container)
    \$ ./scripts/run_bench.sh > tally_results/b1.log 2>&1
    (On host machine)
    \$ sudo nvidia-smi -i 0 -c EXCLUSIVE_PROCESS
    (In docker container)
    \$ ./scripts/run_bench.sh > tally_results/b2.log 2>&1
A.6. Evaluation and expected results
  • 1.
    Parse results into CSV files.
    (On host or in container)
    \$ python3 ./scripts/parse_results.py
  • 2.
    Generate plots.
    (On host or in container)
    \$ python3 ./scripts/plot_results_micro.py
  • 3.
    Plots similar to Figure 5 and Figure 7(b) can be found in the tally_results directory.
A.7. Experiment customization
To run the full benchmark, pass the --run-full-benchmark flag to run_bench.sh. Use the scripts/plot_results.py script to generate the plots. Additional scripts for running other experiments and producing figures included in the paper are available in the scripts directory.